# On GRPO Collapse in Search-R1: The Lazy Likelihood-Displacement Death Spiral

### v1

@@ 6.2 Ablation Study and Analysis
Impact of response-level gating. Applying response-level gating reduces the strength of the regularization applied during training, leading to measurable improvements on multi-hop QA tasks. As shown in Tab.˜1, although the average performance increases modestly by 0.2%, the method yields a notable 1.6% gain on the Bamboogle dataset, highlighting its effectiveness in enhancing multi-hop reasoning.
NQ vs. NQ+Hotpot. Models trained only on NQ achieve strong performance on open-domain, single-hop QA but show limited ability to generalize to multi-hop settings. Moving to NQ+Hotpot consistently improves multi-step reasoning: for example, on Qwen2.5-3B-Base, the GRPO baseline increases from 0.303 (NQ-only) to 0.312 (NQ+Hotpot), and LLDS increases from 0.323 to 0.360 under the same shift. A similar trend appears for Qwen2.5-7B-Base, where LLDS rises from 0.417 (NQ-only) to 0.462 (NQ+Hotpot). These gains show that adding multi-hop HotpotQA, whose questions require multi-turn reasoning to answer correctly and receive reward, encourages the model to retrieve, integrate, and reason over multiple pieces of evidence more effectively than training on NQ alone.
@@
Impact of masking answer (MA). We apply masking answer (MA) specifically to model variants where vanilla GRPO and LLDS degenerate into issuing only a single search call, limiting their ability to collect sufficient evidence for multi-hop reasoning. MA disables LLDS regularization on the final answer tokens, encouraging the model to perform additional search steps or intermediate reasoning. As shown in Tab.˜1, adding MA on top of LLDS yields substantial improvements for these one-search–limited models. LLDS-MA increases the Qwen2.5-3B-Base average score in the NQ+Hotpot setting from 0.360 (LLDS) to 0.430, and similarly improves the NQ-only Instruct variant from 0.319 to 0.381. These gains demonstrate that MA effectively encourages deeper, multi-step reasoning in scenarios where the underlying GRPO policy underutilizes search actions. More detailed analysis see Sec.˜B.3.
6.3 Effect of LLDS on Training Stability across Models
To verify the universality of the training collapse issue in GRPO and the robustness of our proposed solution, we extended our experiments across different model scales and alignment stages. Fig.˜6 illustrates the training reward curves for Qwen-2.5 3B and 7B models, covering both Base and Instruct versions. As observed in the blue curves, the baseline GRPO consistently suffers from catastrophic collapse within the first 300 steps, regardless of the model size (3B vs. 7B) or type (Base vs. Instruct). This confirms that the instability is an inherent characteristic of the algorithm rather than an artifact of a specific model configuration. In contrast, the integration of LLDS (red curves) effectively mitigates this issue. In all four scenarios, the models trained with LLDS maintain a steady upward trajectory in rewards, successfully bypassing the collapse points that affect the baseline. This demonstrates that LLDS serves as a generalized and effective regularizer for stabilizing GRPO training.
(a) Qwen-2.5-3B-Base
(b) Qwen-2.5-3B-Instruct
(c) Qwen-2.5-7B-Base
(d) Qwen-2.5-7B-Instruct
Figure 6: Comparisons of training reward curves between baseline GRPO (blue) and GRPO + LLDS (red). The baseline consistently collapses (reward drops to zero) across all model sizes, while LLDS stabilizes training and sustains high rewards.
7 More Discussion and Guideline
Understand why tool-integrated GRPO is uniquely vulnerable to LLD. Unlike free-form RL on text, tool-augmented agents introduce structural conditions that magnify likelihood drift. First, tool calls inject inherently out-of-distribution tokens—such as search results, API outputs, or error messages—that differ sharply from pretrained language distributions. These OOD segments raise token-level uncertainty and make GRPO’s relative updates more volatile, accelerating the onset of likelihood displacement. Second, tool-based reasoning unfolds across multiple stages, with early stages (e.g., query formulation) stabilizing more quickly than later ones (e.g., tool-result interpretation). Because GRPO applies a single scalar reward to all tokens, early-stage tokens that share nearly identical prefixes across correct and incorrect trajectories receive conflicting gradient signals. This reward–token misalignment disproportionately harms stable prefixes and amplifies LLD. Recognizing these structural challenges is essential for diagnosing and mitigating instability in tool-integrated RL systems.
Closely monitor likelihood dynamics—reward alone is insufficient. A central insight of our analysis is that likelihood degradation begins long before any visible drop in reward. Across models, reward continues to rise throughout both the early and steady-decay phases, even as the likelihood of correct responses is already drifting downward. Because GRPO updates are governed by likelihood ratios, such early degradation silently amplifies gradients and initiates instability. Monitoring only the reward therefore obscures the onset of LLD. In practice, reliable early-warning signals come from tracking action-level and trajectory-level log-likelihood, inspecting entropy trajectories, which surge sharply once likelihood begins to collapse, and watching for sudden spikes in likelihood ratios that indicate gradient amplification. Visualization of likelihood or entropy curves provides the earliest and most trustworthy indication that training is approaching the unstable regime.
@@ 8 Conclusion
In this work, we firstly identified Lazy Likelihood Displacement as the central cause of instability in GRPO-based, Search-R1–style tool-integrated RL. We showed that LLD emerges early, occurs frequently, and reliably drives training into a self-reinforcing LLD Death Spiral characterized by monotonic likelihood decay, entropy spikes, gradient inflation, and eventual collapse. Our analysis provides the first systematic evidence that GRPO is uniquely vulnerable to likelihood drift in tool-integrated settings. To counteract this failure mode, we proposed $\rm LLDS$, a simple and effective likelihood-preserving regularizer that activates only when likelihood decreases and targets only the offending tokens. Empirically, $\rm LLDS$ stabilizes optimization and yields consistent gains across seven QA benchmarks, with improvements of up to +37.8% on Qwen2.5-3B and +32.0% on Qwen2.5-7B. These findings highlight likelihood control as a key mechanism for reliable RL-based tool use and offer a promising foundation for more robust and scalable training methods in future agentic LLMs.
@@ Appendix A Theorem and Proof
Theorem A.1 (Trajectory-Level LLD in Tool-Integrated GRPO)
@@
    Embedding similarity: When incorrect responses are similar to correct ones, their prefix representations exhibit large inner products, amplifying the negative contribution. This high representational overlap means the model struggles to disentangle correct from incorrect continuations, causing negative examples to push gradients in harmful directions and make the model unconfident.
@@
Theorem A.2 (Action-level)
@@
$\alpha^{-}_{(i,t,k),(j,t^{\prime},k^{\prime})}=\left\langle\mathbf{e}_{y^{+}_{i,t,k}}-\pi_{\theta(s)}(\cdot\mid x,y^{+}_{i,<t},o^{+}_{i,<t},y^{+}_{i,t,<k}),\mathbf{e}_{y^{-}_{j,t^{\prime},k^{\prime}}}-\pi_{\theta(s)}(\cdot\mid x,y^{-}_{j,<t^{\prime}},o^{-}_{j,<t^{\prime}},y^{-}_{j,t^{\prime},<k^{\prime}})\right\rangle$ and
@@ Appendix B Additional Analysis
B.1 Training Instability Prior to Collapse
Figure 7: Demonstration of unstability prior collapse.
While the final collapse is indeed triggered by large gradients from negative responses, substantial instability and a noticeable performance decline had already emerged earlier in training. As shown in Fig.˜7, the brown vertical line marks the first sharp drop in model reward (green curve), during which the likelihood of correct responses also decreases significantly, even though the gradient norm remains relatively small (around 2). As training progresses, the LD issue intensifies and the low confidence responses lead to gradient norm steadily grows, ultimately causing the model to collapse. These observations suggest that instability is present well before the collapse point and therefore warrants close attention.
B.2 Impact of Regularization Strength
To examine the effect of regularization strength, we ablate $\lambda\in{0,0.01,0.1}$ and plot the corresponding training reward dynamics in Fig.˜8. As shown, without regularization ($\lambda=0$), the model collapses around step 200. A small regularization value ($\lambda=0.01$, orange curve) delays but does not prevent collapse, which occurs around step 220. In contrast, a stronger regularization ($\lambda=0.1$) stabilizes training entirely, allowing the model to continue smoothly without collapse. We truncate the plot at step 250 for clarity, although training can proceed well beyond this point.
Figure 8: Impact of regularization strength on Qwen2.5-7B-base.
B.3 Unlocking Multi-step Reasoning via Answer Masking
To further explore the potential of LLDS in promoting complex reasoning behaviors, we conducted an ablation study on the Qwen-2.5-3B-Base model, focusing on the regularization scope. Specifically, we mask out the final answer tokens from the calculation of the $L_{\text{LLDS}}$ term (denoted as "Mask Answer").
It is important to note that Qwen-2.5-3B-Base inherently lacks the capability for multi-turn tool calling. As illustrated in Fig.˜9, this limitation is evident in the standard training setups: the baseline GRPO (blue curve) rapidly suffers from model collapse, with the number of valid searches dropping to zero. While the standard GRPO + LLDS (red curve) successfully stabilizes the training and prevents collapse, it fails to elicit any multi-step behavior, with the search frequency remaining stagnant at exactly 1.0.
However, the "GRPO + LLDS-MA" variant (green curve) demonstrates a distinct emergent behavior. By removing the regularization constraint on the final answer tokens, we observe a significant increase in the number of valid searches per question, rising well above 2.0. This indicates that relaxing the penalty on the answer tokens effectively unlocks the model’s latent potential to engage in multi-step reasoning and utilize external tools—capabilities that standard methods failed to activate in the base model.
Figure 9: Evolution of the number of valid searches per question on Qwen-2.5-3B-Base. The base model inherently lacks multi-turn tool calling capabilities. While the baseline (blue) collapses and standard LLDS (red) stabilizes at a single search step, LLDS-MA (green) successfully unlocks the model’s potential for multi-step reasoning, driving the number of valid searches significantly above 1.0.
B.4 Qualitative Examples of Tool-Integrated Reasoning

### v2

@@ 6.2 Ablation Study and Analysis
Impact of Regularization Strength. We ablate $\lambda\in\{0,0.01,0.1,0.2\}$ and plot the corresponding training reward dynamics in Figure˜8. As shown, Vanilla GRPO ($\lambda=0$) collapses around step 200. A small regularization value ($\lambda=0.01$, orange curve) delays but does not prevent collapse, which occurs around step 220. In contrast, a stronger regularization ($\lambda=0.2$) stabilizes training entirely, allowing the model to continue smoothly without collapse. We truncate the plot at step 500 for clarity, although training can proceed well beyond this point.
Figure 8: Impact of regularization strength in training Qwen2.5-3B-Instruct on NQ and HotpotQA.
@@
Impact of masking answer (MA). We apply MA only to model variants where vanilla GRPO degenerate into issuing only a single search call. MA reduces LLDS regularization on the final answer tokens, encouraging the model to perform additional search steps or intermediate reasoning. As shown in Table˜1, LLDS-MA increases the Qwen2.5-3B-Base average score in the NQ+Hotpot setting from 0.360 (LLDS) to 0.430. These gains demonstrate that MA effectively encourages deeper, multi-step reasoning in scenarios where the underlying GRPO policy underutilizes search actions. More detailed analysis see Section˜B.4.4.
Effect of LLDS on Training Stability across Models To assess the universality of GRPO collapse and the robustness of our method, we evaluate across model scales (3B and 7B) and alignment stages (Base and Instruct). As shown in Figure˜7, vanilla GRPO consistently collapses within the first 300 steps regardless of model size or variant. In contrast, integrating $\rm LLDS$ (red line) reliably stabilizes training and maintains steady reward growth across all four settings, successfully avoiding the collapse issue. These results demonstrate that $\rm LLDS$ serves as a general and effective regularizer for stabilizing GRPO training.
@@ 7 Conclusion
In this work, we identify LLD as the primary cause of failure in GRPO-based, search-based tool-integrated reinforcement learning, and show that it emerges early and evolves into a self-reinforcing failure mode characterized by likelihood decay, entropy inflation, and gradient explosion. Our analysis reveals that tool integration fundamentally reshapes group-relative optimization dynamics by coupling likelihood updates across trajectories, leading to a class of instability unique to agentic, tool-using LLMs. To counteract this failure mode, we propose $\rm LLDS$, a simple and effective likelihood-preserving regularizer that activates only when likelihood decreases and targets only the offending tokens, thereby stabilizing training and delivering consistent gains across seven QA benchmarks. More broadly, we highlight likelihood dynamics as a principled signal for early failure detection, and suggest that extending $\rm LLDS$ to other tools and agentic settings is a promising direction for future work.
@@ Appendix A Theorem and Proof
Theorem A.1 (Trajectory-Level LLD in Tool-Integrated GRPO).
@@
    Embedding similarity: When incorrect responses are similar to correct ones, their representations exhibit large inner products, amplifying the negative contribution. This high representational overlap means the model struggles to disentangle correct from incorrect continuations, causing negative examples to push gradients in harmful directions and make the model unconfident.
@@
Theorem A.2 (Action-level).
@@
$\alpha^{-}_{(i,t,k),(j,t^{\prime},k^{\prime})}=\left\langle\mathbf{e}_{\bm{y}^{+}_{i,t,k}}-\pi_{\theta(s)}(\cdot\mid\bm{x},\bm{y}^{+}_{i,<t},\bm{o}^{+}_{i,<t},\bm{y}^{+}_{i,t,<k}),\mathbf{e}_{\bm{y}^{-}_{j,t^{\prime},k^{\prime}}}-\pi_{\theta(s)}(\cdot\mid\bm{x},\bm{y}^{-}_{j,<t^{\prime}},\bm{o}^{-}_{j,<t^{\prime}},\bm{y}^{-}_{j,t^{\prime},<k^{\prime}})\right\rangle$ and
@@
A.2 Formal Definition of the LLD Death Spiral
Definition A.3 (LLD Death Spiral).
[…]
We define the token-level likelihood change at iteration $s$ as
[…]
and the trajectory-level likelihood change as
$\Delta^{(s)}_{i}=\sum_{t=0}^{T}\sum_{k=1}^{|y^{+}_{i,t}|}\Delta^{(s)}_{i,t,k}$.
The training dynamics are said to enter an LLD Death Spiral at iteration $s_{0}$ if there exist constants $\varepsilon<0$ and $\delta>0$ such that, for all $s\geq s_{0}$, the following conditions hold:
  1. 1.
    Persistent likelihood decay. The expected likelihood of correct trajectories decreases monotonically:
    $\mathbb{E}_{i}\!\left[\Delta^{(s)}_{i}\right]\leq\varepsilon.$
  2. 2.
    Confidence erosion under response similarity. As likelihood decays, the model assigns no higher probability mass to incorrect trajectories,
    $\mathbb{E}_{\hat{\bm{y}}^{-}\sim\pi_{\theta_{s}}(\cdot\mid\bm{x}_{i})}\!\left[\pi_{\theta_{s}}(\hat{\bm{y}}^{-}\mid\bm{x}_{i})\right]\;\leq\;\mathbb{E}_{\hat{\bm{y}}^{-}\sim\pi_{\theta_{s-1}}(\cdot\mid\bm{x}_{i})}\!\left[\pi_{\theta_{s-1}}(\hat{\bm{y}}^{-}\mid\bm{x}_{i})\right],$
    while incorrect and correct trajectories remain representationally similar:
    $\mathbb{E}_{\hat{\bm{y}}^{-},\hat{\bm{y}}^{+}\sim\pi_{\theta_{s}}(\cdot\mid\bm{x}_{i})}\!\left[\mathrm{sim}\!\left(\bm{h}(\bm{x}_{i},\hat{\bm{y}}^{-}),\,\bm{h}(\bm{x}_{i},\hat{\bm{y}}^{+})\right)\right]\geq c,$
    for some constant $c>0$  [^1] In GRPO, correct and incorrect responses often exhibit similarity due to shared conditioning on the same query; see Figure 5.. Here $\bm{h}(\cdot)$ denotes the policy’s hidden representation.
  3. 3.
    Self-amplifying decay. Lower response likelihood causes likelihood decay to become self-amplifying over training iterations.:
    $\mathbb{E}_{i}\!\left[\Delta^{(s+1)}_{i}\right]\;<\;\mathbb{E}_{i}\!\left[\Delta^{(s)}_{i}\right$],
    i.e., the magnitude of expected likelihood decay strictly increases over training.
Lowlikelihood of Incorrect responses. We demonstrate the likelihood of correct and incorrect responses in Figure˜9(a). As shown, after step 80, the likelihood of incorrect responses is consistently lower than that of correct responses. This provides evidence that incorrect trajectories lie in a low-probability regime of the policy, which causes their token-level prediction errors to be assigned disproportionately large weights in the GRPO gradient. As a result, these low-likelihood incorrect responses exert an outsized negative influence on similar correct trajectories, suppressing the likelihood of correct actions and accelerating Lazy Likelihood Displacement. This observation is consistent with the theoretical analysis that low-probability negative samples amplify gradient interference and contribute directly to the onset of the LLD death spiral.
@@ Appendix B Additional Analysis
B.1 Detail Related Work
B.1.1 Tool-Integrated Reasoning and Agentic LLMs.
Tool use has emerged as a powerful paradigm for equipping LLMs with adaptive reasoning capabilities. Early approaches relied on prompt-based orchestration (Lu et al., 2023; Shen et al., 2023) or multi-agent delegation frameworks to invoke tools without explicit training. Instruction-tuned models (Gou et al., 2023; Qin et al., 2023) later introduced structured tool-calling behaviors via supervised learning, but these systems remained largely static and were constrained to single-turn interactions. More recent work demonstrates that reinforcement learning can substantially enhance tool integration by enabling models to learn tool-usage policies through environmental feedback and task success. Notable systems such as RETool (Feng et al., 2025) and VERL-Tool (Jiang et al., 2025) support multi-step reasoning through dynamic tool use, self-verification, and error correction. This transition from static instruction-following to feedback-driven optimization has proven effective across a range of domains, including mathematical problem solving with code execution (Xue et al., 2025), open-domain question answering with retrieval (Jin et al., 2025), SQL generation from natural language (Jiang et al., 2025), and multi-modal visual reasoning (Gao et al., 2024).
B.1.2 Training Collapse in Tool-Integrated GRPO.
GRPO (Guo et al., 2025) has gained popularity in reinforcement learning due to its value-free, outcome-driven formulation. However, applying GRPO to train LLMs for multi-turn tool-integrated reasoning remains highly unstable (Jin et al., 2025). When initialized from base models using only verifiable rewards (Zero RL), GRPO-trained models often experience catastrophic failure, marked by abrupt reward collapse (Xue et al., 2025; Jin et al., 2025). In contrast, Proximal Policy Optimization (PPO) (Schulman et al., 2017) tends to exhibit more stable behavior under comparable settings (Jin et al., 2025).
These collapse dynamics have been consistently observed across various systems, including Search-R1 (Jin et al., 2025), SimpleTIR (Xue et al., 2025), and ZeroSearch (Sun et al., 2025). While prior work has extensively documented these failures empirically, principled explanations remain limited. Among existing studies, (Xue et al., 2025) attributes the instability to low-likelihood incorrect responses that induce excessively large importance weights. Separately, (Liu et al., 2025) argues that off-policy accelerated inference introduces training–inference mismatch that exacerbates collapse. However, neither explanation accounts for why structurally similar algorithms such as PPO do not suffer comparable failures, nor do they clarify the deeper structural origin of these low-likelihood trajectories. Moreover, we observe that reward degradation often precedes the final collapse, suggesting a more fundamental optimization pathology rather than a purely stochastic failure. In this work, we identify LLD as the core mechanism underlying GRPO’s failure mode in multi-turn TIRL.
To mitigate this issue, recent works introduce turn-level reward shaping for improved credit assignment (Zheng et al., 2025b; Zhang et al., 2025; Ji et al., 2025). StepSearch (Zheng et al., 2025b) constructs turn-wise rewards using ground-truth supporting documents, which are often unavailable in realistic settings. CriticSearch (Zhang et al., 2025) relies on an external LLM to judge each action, incurring additional computational cost and dependence on auxiliary model capability. Moreover, dense intermediate rewards could be susceptible to reward hacking (Guo et al., 2025). Recently, TreeGRPO (Ji et al., 2025), provide dense reward using tree-based rollout, but increases rollout complexity and computational overhead. In contrast, our approach shows that simple rule-based outcome rewards alone can achieve superior performance once GRPO training is properly stabilized.
B.2 Additional Application Details
B.2.1 Hyper-parameters
Following Search-R1 (Jin et al., 2025), we implement our system using vLLM with a tensor parallelism degree of 1 and a GPU memory utilization ratio of 0.6. Rollout sampling is performed with temperature 1.0 and top-$p$ set to 1.0. We set the KL-divergence regularization coefficient to 0.001 and the clipping ratio to 0.2. For GRPO training, the policy LLM is optimized with a learning rate of $1\times 10^{-6}$, and five responses are sampled per prompt. Training is conducted on four H100 GPUs with a total batch size of 512, a mini-batch size of 256, and a micro-batch size of 64. The maximum sequence length is 4096 tokens, including up to 500 tokens for the generated response and 500 tokens for retrieved content. In addition, we reduce the maximum number of turns from 4 to 3 for efficiency. Although this change favors Search-R1, our method still outperforms it under this setting. During evaluation, the maximum action budget is set to 4, and the top-3 passages are retrieved by default.We set $\lambda=0.2$ for all models, except for Qwen-2.5-7B-Instruct where $\lambda=0.3$. For $\rm LLDS$-MA on Qwen-2.5-3B-Base, we linearly increase $\beta$ from 0 to $1$ according to $\beta=\min\!\left(\frac{V_{s}}{V_{B}},\,1\right),$ where $V_{s}$ denotes the average number of valid searches at time $s$ and $V_{B}$ is a baseline value fixed to $2$.
B.2.2 Baseline Methods
Direct Inference performs single-pass answer generation without external retrieval or iterative reasoning, serving as a minimal baseline to measure the intrinsic capability of the base LLM.
Search-o1 (Li et al., 2025) introduces structured search planning with predefined search-action templates, enabling more controlled interaction with the retriever. However, it remains inference-only and does not learn to optimize search policies from task feedback.
Retrieval-Augmented Generation (RAG) (Lewis et al., 2020) retrieves a fixed set of documents once at the beginning and conditions answer generation on the retrieved context. RAG improves factual grounding but lacks iterative refinement, making it less effective for multi-hop reasoning where intermediate queries must adapt to evolving reasoning states.
R1 (Guo et al., 2025) is an RL-based fine-tuning method that optimizes reasoning trajectories using outcome-level rewards, but operates without a search engine. As a result, R1 primarily improves reasoning structure while remaining constrained by the model’s internal knowledge.
Rejection Sampling (Ahn et al., 2024) generates multiple candidate reasoning trajectories and retains only those leading to correct answers for supervised fine-tuning.
Search-R1-PPO / Search-R1-GRPO (Jin et al., 2025) extend R1 by incorporating a search engine during rollout and optimizing search-aware reasoning trajectories via PPO or GRPO.
ZeroSearch (Sun et al., 2025) trains search-aware LLMs via reinforcement learning without interacting with real search engines, by replacing retrieval with a simulated LLM-based search model under a curriculum strategy.
StepSearch (Zheng et al., 2025b) introduces turn-level reward shaping by assigning each search step an information-gain reward computed from similarity to ground-truth supporting documents, minus a redundancy penalty. This provides dense supervision but requires curated step-level document annotations that are often unavailable in realistic settings. We use the official checkpoints to evaluate.
CriticSearch (Zhang et al., 2025) uses a frozen external LLM to retrospectively judge each search action as good or bad based on its contribution to the final answer, converting these judgments into turn-level rewards. While general and effective, it introduces additional computational cost, depends on the capability of the auxiliary judge model, and increases exposure to reward hacking.
TreeGRPO (Ji et al., 2025) samples branching trajectories and propagates outcome rewards backward through the tree to produce step-level advantages via relative comparisons among sibling branches. This yields denser credit assignment but significantly increases rollout complexity and computational overhead.
B.3 LLDS Variants
LLDS: Token-Level Likelihood Preservation.
One intial variant is the token-level penalty: within each preserved trajectory $\bm{y}_{i}$ and each action (turn) $\bm{y}_{i,t}=(y_{i,t,1},\dots,y_{i,t,|\bm{y}_{i,t}|})$, only likelihood-reducing tokens contribute to the loss. Then
$L_{\rm LLDS(Token)}$ $=\frac{1}{\sum_{\bm{y}_{i}\in\mathcal{Y}_{\rm pre}}\sum_{t=0}^{T_{i}}|\bm{y}_{i,t}|}\sum_{\bm{y}_{i}\in\mathcal{Y}_{\rm pre}}\sum_{t=0}^{T_{i}}\sum_{k=1}^{|\bm{y}_{i,t}|}\underbrace{\max(0,\Delta_{i,t,k})}_{\text{likelihood-reducing tokens}}.$ (9)
LLDS: Response-Level Gating.
To avoid penalizing globally improving actions, $\rm LLDS$ introduces a response-level gating mechanism: the penalty activates only when the total likelihood of the response decreases. The LLDS loss is:
$L_{\rm LLDS(response)}$ $=\frac{1}{\sum_{\bm{y}_{i}\in\mathcal{Y}_{\rm pre}}\sum_{t=0}^{T_{i}}|\bm{y}_{i,t}|}\sum_{\bm{y}_{i}\in\mathcal{Y}_{\rm pre}}\mathbbm{1}\!\left[\sum_{t=0}^{T_{i}}\sum_{k=1}^{|\bm{y}_{i,t}|}\Delta_{i,t,k}>0\right]\sum_{t=0}^{T_{i}}\sum_{k=1}^{|\bm{y}_{i,t}|}\underbrace{\max(0,\Delta_{i,t,k})}_{\text{likelihood-reducing tokens}}.$ (10)
However, a mismatch can arise between action-level and response-level likelihood changes: an individual action may suffer from LLD even when the overall response improves, or conversely, the response may exhibit LLD while only a subset of actions deteriorate. As a result, response-level gating can lead to either over-penalization or under-penalization at the action granularity. To address this issue, we adopt an action-level variant of $\rm LLDS$, defined in Equation˜4. Unless stated otherwise, $\rm LLDS$ refers to the action-level gating formulation throughout this paper.
B.4 Additional Results
B.4.1 Performance Comparison Among LLDS Variants
We conduct an ablation study to compare different variants of $\rm LLDS$, including token-level, response-level, and action-level gating. The results are summarized in Table˜3. We first evaluate the three variants under the NQ-only training setting on Qwen2.5-3B-Base. The token-level variant yields the lowest performance among the three, indicating that penalizing all likelihood-reducing tokens, even within globally improving actions or responses, introduces overly strong regularization that interferes with policy optimization. We further compare the response-level and action-level variants under the more challenging NQ+Hotpot training setting on both Qwen2.5-3B-Base and Qwen2.5-3B-Instruct. In all cases, the action-level variant achieves the best performance across both general QA and multi-hop QA benchmarks, improving the overall average score by a consistent margin (e.g., from 0.353 to 0.360 on Qwen2.5-3B-Base and from 0.419 to 0.441 on Qwen2.5-3B-Instruct). While all variants outperform vanilla GRPO, action-level gating achieves the strongest empirical performance.
Methods General QA Gen-Avg Multi-Hop QA MH-Avg Avg
NQ TriviaQA PopQA HotpotQA 2Wiki Musique Bamboogle
Qwen2.5-3b-Base
NQ-Only
+LLDS (token) 0.462 0.609 0.464 0.512 0.286 0.248 0.063 0.113 0.178 0.321
+LLDS (response) 0.478 0.605 0.449 0.511 0.288 0.250 0.062 0.129 0.182 0.323
+LLDS 0.491 0.615 0.450 0.519 0.306 0.275 0.067 0.121 0.192 0.332
NQ+Hotpot
+LLDS (response) 0.480 0.631 0.455 0.522 0.352 0.311 0.087 0.153 0.226 0.353
+LLDS 0.479 0.627 0.453 0.520 0.345 0.323 0.082 0.210 0.240 0.360
Qwen2.5-3b-Ins
NQ-Only
+LLDS (token) 0.469 0.609 0.457 0.512 0.291 0.250 0.057 0.097 0.174 0.319
+LLDS (response) 0.490 0.610 0.450 0.517 0.294 0.193 0.074 0.121 0.171 0.319
+LLDS 0.485 0.614 0.455 0.517 0.299 0.266 0.062 0.129 0.189 0.330
NQ+Hotpot
+LLDS (response) 0.462 0.622 0.460 0.515 0.432 0.383 0.180 0.395 0.347 0.419
+LLDS 0.462 0.627 0.468 0.519 0.443 0.447 0.197 0.444 0.383 0.441 (+31.3%)
Search-R1-GSPO 0.376 0.561 0.374 0.437 0.332 0.329 0.098 0.250 0.252 0.331
+LLDS (response) 0.484 0.637 0.475 0.532 0.452 0.448 0.212 0.387 0.375 0.442(+33.5%)
+LLDS 0.489 0.640 0.483 0.537 0.466 0.456 0.215 0.403 0.385 0.451(+36.3%)
Table 3: Results of different $\rm LLDS$ variants on General QA and Multi-Hop QA datasets.
B.4.2 Results of training with NQ dataset only.
In this section, we report results using training on the NQ dataset only. Models trained solely on NQ achieve strong performance on open-domain single-hop QA, but exhibit limited generalization to multi-hop reasoning. Incorporating HotpotQA (NQ+Hotpot) consistently enhances multi-step reasoning ability. As shown in Table˜3, for Qwen2.5-3B-Base, the GRPO baseline improves from 0.303 (NQ-only) to 0.312 (NQ+Hotpot), while $\rm LLDS$ (response) increases from 0.323 to 0.353, and $\rm LLDS$ further improves from 0.332 to 0.360 under the same setting. These gains indicate that adding multi-hop supervision from HotpotQA, where questions require multi-step reasoning to obtain correct rewards, encourages the model to retrieve, integrate, and reason over multiple pieces of evidence more effectively than training on NQ alone.
B.4.3 Additional Results on Llama.
In this section, we report results on Llama3.2-3B-Base. As shown in Table˜4, vanilla GRPO fails to learn tool use, collapsing into a “reason-and-answer” policy. This failure is reflected in its weak performance, achieving only 0.212 overall average and a particularly low 0.142 MH-Avg on multi-hop benchmarks. By incorporating $\rm LLDS$, the model is successfully stabilized and learns to invoke tools, yielding substantial gains across both general and multi-hop QA. Specifically, Gen-Avg improves from 0.305 to 0.532, MH-Avg rises from 0.142 to 0.255, and the overall average increases from 0.212 to 0.374. However, the learned tool behavior remains being single-turn usage. Finally, MA is applied on top of LLDS, the model is further encouraged to engage in multi-turn tool use, unlocking additional improvements on multi-hop reasoning. The MH-Avg increases from 0.255 to 0.343, while maintaining strong general QA performance (Gen-Avg = 0.538), leading to the best overall average of 0.427.
Methods General QA Gen-Avg Multi-Hop QA MH-Avg Avg
NQ TriviaQA PopQA HotpotQA 2Wiki Musique Bamboogle
GRPO 0.233 0.461 0.220 0.305 0.196 0.232 0.044 0.097 0.142 0.212
GRPO-LLDS 0.486 0.643 0.467 0.532 0.366 0.298 0.113 0.242 0.255 0.374
GRPO-LLDS-MA 0.493 0.640 0.480 0.538 0.449 0.462 0.164 0.298 0.343 0.427
Table 4: Results on General QA and Multi-Hop QA datasets for Llama3.2-3b-Base.
(a) Likelihood of correct and incorrect responses.
(b) Evolution of the number of valid searches per question on Qwen-2.5-3B-Base.
Figure 9: Analysis of Likelihood dynamics and tool usage behaviors. Left: Likelihood comparison between correct and incorrect responses using NQ+Hotpot dataset. Right: Comparison of valid search steps across different training strategies.
B.4.4 Encouraging Multi-step Reasoning via Answer Masking
As shown in Figure˜9(b), the Qwen2.5-3B-Base model lacks inherent multi-turn tool-use capability: vanilla GRPO (blue) rapidly collapses, with search frequency fixed at 1.0, while GRPO+LLDS (red) stabilizes training but remains a single-turn search. In contrast, GRPO+LLDS-MA (green), which reduces regularization on answer tokens, exhibits clear emergent behavior, increasing valid searches per query beyond 2.0. This indicates that relaxing answer-token penalties unlocks latent multi-step reasoning and tool-use capabilities absent under standard training.
B.5 Behavioral Characteristics